# KAN KAN Buff Signed Graph Neural Networks?

### v1

@@ 4.2 Parameters
Each KAN layer consists of a grid of splines. The degree of each spline is the spline order, and the number of splines for each function is known as grid size. For the implementation of B-splines and other variants of KAN, we rely on the publicly available implementations in [Blealtan, 2024] and [Yin et al., 2024], respectively. The parameters used in the experiments are: layers = [32,32], weight_decay = $10^{-5}$, learning_rate = 0.001, lamb = 1.0, epochs = 1000, seed = 42, reduction_iterations = 10, reduction_dimensions = 15, spectral_features = True, norm = True, norm_embed = True, grid_size = 5, spline_order = 3, scale_noise = 0.1, scale_base = 1.0, scale_spline = 1.0, base_activation = torch.nn.SiLU, grid_eps = 0.02, grid_range= [-1, 1], Kmeans++ is used for clustering and with default parameters. Each experiment is repeated 10 times and averaged. The test size split for each positive and negative for link sign prediction (using multinomial logistic regression) is 0.2. t-SNE parameters are random_state = 0, n_iter = 1000, metric = ’cosine’. Preprocessing of the signed network includes Removing duplicates and keeping a first edge, treating neutral edges as positive, dropping self-loops, reindex nodes from 0 to $n-1$ where $n$ is the number of nodes in the signed network. A tanh activation function is applied after passing the embedding to the KAN layer.
@@ 4.3 Evaluation Metrics
where $pos_{within}$ and $pos_{between}$ are the number of positive edges within and between clusters, respectively. $neg_{within}$ and $neg_{between}$ are the negative edges within and between clusters. We rely on the AUC score that summarizes the GNN’s effectiveness across all possible classification thresholds to evaluate the link sign prediction performance. In addition, we use the F1 score for this as well, which is as follows:
@@ 4.4 Signed Graphs
4.5 Tasks
4.5.1 Community Detection
  • Research Question: Does incorporating the KAN layer into the signed graph convolutional network improve the clustering quality when using Kmeans++ on the embeddings?
@@ 4.5.1 Community Detection
    Table 2: Average $pos\_{in}$ and $neg\_{out}$ over 10 runs for SGCN and KASGCN node embeddings using Kmeans++ clustering ($K=5$).
@@
    Table 3: Average $pos\_{in}$ and $neg\_{out}$ over 10 runs for SGCN and KASGCN node embeddings using Kmeans++ clustering ($K=10$).
@@
    Table 4: Average $pos\_{in}$ and $neg\_{out}$ over 10 runs for SGCN and KASGCN node embeddings using Kmeans++ clustering ($K=15$).
@@
  • Observations: Improvement in the clustering quality highly depends on the choice of K. For relatively higher values of K, Kmeans++ using embeddings from KASGCN seems to do slightly better. In comparison, for relatively lower values of K, Kmeans++ using embeddings from SGCN seems to do slightly better. Overall, the difference in the clustering quality between SGCN and KASGCN is negligible or meager except in PPI, where KASGCN outperforms SGCN regardless of K.
4.5.2 Link Sign Prediction
  • Research Question: Do the embeddings generated from integrating the KAN layer into SGCN increase the predictive power of logistic regression for sign classification?
@@ 4.6 Efficiency Examination
  • Research Question: How efficient is KASGCN compared to SGCN?
@@
  • Observations: As expected, due to the spline function, KAGCN takes a significantly longer time to train and generate embeddings. Increasing the aggregation layers impacts the execution time for KASGCN but not SGCN.
@@ 4.7 Embeddings Comparison
  • Research Question: Given that the observed improvements are relatively modest, how similar do SGCN and KASGCN produce the embeddings? Are these slight enhancements possible due to the similarity in the embeddings generated by the two models?
@@
  • Results: The embeddings for SGCN (Red) and KASGCN (Blue) are visualized using t-SNE for BitcoinAlpha, BitcoinOTC, and WikiElec in Figure 2. The average cosine similarity between the embeddings of SGCN and KASGCN for these three signed graphs is computed to be -0.02, 0.01, and 0.072, respectively.
@@
4.8 KAN Variants Trial
  • Research Question: Does integrating other KAN variants into SGCN perform just as well as integrating the original KAN layer?
@@ 4.8 KAN Variants Trial
    Table 6: Average $pos\_{in}$ and $neg\_{out}$ over 10 runs for KASGCN, FourierKASGCN, LaplaceKASGCN, and WaveletKASGCN node embeddings using Kmeans++ clustering ($K=5$).
@@
    Table 7: Average $pos\_{in}$ and $neg\_{out}$ over 10 runs for KASGCN, FourierKASGCN, LaplaceKASGCN, and WaveletKASGCN node embeddings using Kmeans++ clustering ($K=10$).
@@
    Table 8: Average $pos\_{in}$ and $neg\_{out}$ over 10 runs for KASGCN, FourierKASGCN, LaplaceKASGCN, and WaveletKASGCN node embeddings using Kmeans++ clustering ($K=15$).
@@
    Table 9: Average AUC and F1 score over 10 runs for KASGCN, FourierKASGCN, LaplaceKASGCN, and WaveletKASGCN node embeddings using Multinomial Logistic Regression with random 0.2 test size split.
@@
5 Summary
In this study, we integrated the Kolmogorov-Arnold Neural Network into the Signed Graph Convolutional Network to investigate its effectiveness in several tasks, including community detection and link sign prediction in signed networks. Overall, our empirical evaluations demonstrated that KAN-enhanced SGCNs achieve comparable performance compared to vanilla SGCNs, considering the variability in the results indicated by the relatively large standard deviation, and it is context-dependent (signed graph, parameters..etc.). Our study provides a foundation and paves the way for further research concerning the applicability of KANs in various domains.

### v3

@@ 2.2 Parameters
Each KAN layer consists of a grid of splines. The degree of each spline is the spline order, and the number of splines for each function is known as grid size. For the implementation of B-splines and other variants of KAN, we rely on the publicly available implementations in [Blealtan, 2024] and [Yin et al., 2024], respectively. The parameters used in the experiments are: layers = [32,32], weight_decay = $10^{-5}$, learning_rate = 0.001, lamb = 1.0, epochs = 1000, seed = 42, reduction_iterations = 10, reduction_dimensions = 15, spectral_features = True, norm = True, norm_embed = True, grid_size = 5, spline_order = 3, scale_noise = 0.1, scale_base = 1.0, scale_spline = 1.0, base_activation = torch.nn.SiLU, grid_eps = 0.02, grid_range= [-1, 1], Kmeans++ is used for clustering and with default parameters. Each experiment is repeated 10 times and averaged. The test size split for each positive and negative for link sign prediction (using multinomial logistic regression) is 0.2. t-SNE parameters are random_state = 0, n_iter = 1000, metric = ’cosine’. Pre-processing of the signed network includes Removing duplicates and keeping a first edge, treating neutral edges as positive, dropping self-loops, reindex nodes from 0 to $n-1$ where $n$ is the number of nodes in the signed network. The embedding passes through the KAN layer and the tanh activation function.
@@ 2.3 Evaluation Metrics
where $pos_{within}$ and $pos_{between}$ are the number of positive edges within and between clusters, respectively. $neg_{within}$ and $neg_{between}$ are the negative edges within and between clusters. To measure the overall clustering quality $Q$, we add both to obtain:
$Q=pos_{in}+neg_{out}$ (5)
Moreover, we rely on the AUC score that summarizes the GNN’s effectiveness across all possible classification thresholds to evaluate the link sign prediction performance. In addition, we use the F1 score for this as well, which is as follows:
@@ 2.4 Signed Graphs
3 Experimental Results
3.1 Community Detection
@@
    Table 2: Average $pos\_{in}$ and $neg\_{out}$ over 10 runs for SGCN and KASGCN node embeddings using Kmeans++ clustering ($K=5$). Gain is computed using $\frac{100\cdot(Q_{KASGCN}-Q_{SGCN})}{Q_{SGCN}}$.
@@
    Table 3: Average $pos\_{in}$ and $neg\_{out}$ over 10 runs for SGCN and KASGCN node embeddings using Kmeans++ clustering ($K=10$). Gain is computed using $\frac{100\cdot(Q_{KASGCN}-Q_{SGCN})}{Q_{SGCN}}$.
@@
    Table 4: Average $pos\_{in}$ and $neg\_{out}$ over 10 runs for SGCN and KASGCN node embeddings using Kmeans++ clustering ($K=15$). Gain is computed using $\frac{100\cdot(Q_{KASGCN}-Q_{SGCN})}{Q_{SGCN}}$.
@@
  • Observations: Improvement in the clustering quality highly depends on the choice of K. For relatively higher values of K, Kmeans++ using embeddings from KASGCN seems to do slightly better. In comparison, for relatively lower values of K, Kmeans++ using embeddings from SGCN appears to do slightly better. Overall, the difference in the clustering quality between SGCN and KASGCN is negligible or meager except in PPI, where KASGCN outperforms SGCN regardless of K.
@@ 3.3 Efficiency Examination
  • Observations: As expected, due to the spline function, KAGCN takes a significantly longer time to train and generate embeddings. Increasing the aggregation layers increases the execution time for the KASGCN but not for the SGCN.
@@ 3.4 Embeddings Comparison
  • Results: The embeddings for SGCN (Red) and KASGCN (Blue) are visualized using t-SNE for BitcoinAlpha, BitcoinOTC, and WikiElec in Figure 3. The average cosine similarity between the embeddings of SGCN and KASGCN for these three signed graphs is -0.02, 0.01, and 0.072, respectively.
@@
3.5 KAGCN Robustness
@@
    Table 6: Average $pos\_{in}$ and $neg\_{out}$ over 10 runs for KASGCN, FourierKASGCN, LaplaceKASGCN, and WaveletKASGCN node embeddings using Kmeans++ clustering ($K=5$). Numbers in bold indicate that the clustering quality $Q$ is the highest.
@@
    Table 7: Average $pos\_{in}$ and $neg\_{out}$ over 10 runs for KASGCN, FourierKASGCN, LaplaceKASGCN, and WaveletKASGCN node embeddings using Kmeans++ clustering ($K=10$). Numbers in bold indicate that the clustering quality $Q$ is the highest.
@@
    Table 8: Average $pos\_{in}$ and $neg\_{out}$ over 10 runs for KASGCN, FourierKASGCN, LaplaceKASGCN, and WaveletKASGCN node embeddings using Kmeans++ clustering ($K=15$). Numbers in bold indicate that the clustering quality $Q$ is the highest.
@@
    Table 9: Average AUC and F1 score over 10 runs for KASGCN, FourierKASGCN, LaplaceKASGCN, and WaveletKASGCN node embeddings using Multinomial Logistic Regression with random 0.2 test size split. Numbers in bold indicate that the AUC/F1 metric is the highest.
@@
4 Conclusion and Future Work
In this work, we investigated the integration of the Kolmogorov-Arnold Neural Network (KAN) within the Signed Graph Convolutional Network (SGCN) framework, aiming to evaluate its effectiveness in community detection and link sign prediction in signed networks. Our empirical results show that KAN-enhanced SGCNs perform comparably to traditional, unmodified SGCNs. However, the variability in performance, highlighted by relatively large standard deviations, suggests that the effectiveness of this approach is highly context-dependent, influenced by factors such as the specific characteristics of the signed graph and model parameters. Notably, the LaplaceKASGCN variant consistently performs poorly in signed network downstream tasks. The findings in this paper lay the groundwork for future exploration into the broader applicability of KANs in various domains. Specifically for signed graph networks, this research offers valuable insights for further refining the integration of KANs into graph-based learning models, paving the way for more sophisticated and effective analysis in the future.